A snow cone is a tasty treat with flavored ice and a spherical bubble gum ball at the bottom, as shown below:


Snow cone with spherical bubble gum ball at the bottom


The radius of the cone is 1.75 inches, and its height is 3.5 inches. If the diameter of the bubble gum ball is 0.5 inches, what is the closest approximation of the volume of the cone that can be filled with flavored ice?


11.15 in3

11.22 in3

0.07 in3

33.67 in3


1
Expert's answer
2020-07-06T20:37:43-0400

Solution


Volume of the cone is "\frac{1}{3}\cdot(\pi\cdot r^2) \cdot h = \frac{1}{3}\cdot\pi\cdot 1.75^2\cdot3.5 \approx 11.2246" in3

Volume of the sphere is "\frac{4}{3}\cdot\pi\cdot r^3 = \frac{4}{3}\cdot\pi\cdot(\frac{0.5}{2})^3 \approx 0.0654" in3

Volume of the cone without a sphere is "11.2246-0.0654 \approx 11.1592" in3


Answer: 11.15 in3